OpenStudy (anonymous):

the perimeter of a rectangle is 54 m and the area of the rectangle is 72m^2. find the dimensions of the rectangle

OpenStudy (anonymous):

rectangle or square?

OpenStudy (anonymous):

if rectangle A=l*w 54=2l+2w If square A=s^2 54=s*4 Depending on whether square or rectangle solve the system of equation. BUt i bet it is square.

OpenStudy (anonymous):

13.5 sides if a square

OpenStudy (anonymous):

It's not a square, since 13.5^2 != 72 (which was the area) The area of a rectangle is given by w * l, the perimeter is 2w + 2l as timo already said. The area of our rectangle is 72, and the perimeter is 54, so: w * l = 72 2w + 2l = 54 All you have to do is solve the system of equations to get the width and length of the rectangle (hint: solve for w in one equation, and substitute it into the other one)

OpenStudy (anonymous):

i did it wrong i forgot to plug in the area